OpenStudy (anonymous):

How would I graph 8x-9y=-54 8x=9y-54

OpenStudy (shamil98):

First you would put the equation in slope-intercept form. Which is y = mx + b 8x-9y = -54 -9y = -8x - 54 y = 8/9 x + 6

OpenStudy (shamil98):

M represents the slope and b is the y intercept in which you start plotting the rise and run of the graph.

OpenStudy (shamil98):

In this case the rise and run is 8 up and 9 right.

OpenStudy (anonymous):

& Both lines are the same correct?

OpenStudy (shamil98):

Yes, all the second equation is the y variable moved to the other side if you undo that and put in slope-intercept form the result is still the same.

OpenStudy (anonymous):

Thank you

OpenStudy (shamil98):

No problem.